OpenStudy (aravindg):

wel its too simple but can anyone giv me diff forms of using (a+b)^3,a^3+b^3 etc

OpenStudy (anonymous):

\[(a+b)^3=a^3+3a^2b+3ab^2+b^3\] \[a^3+b^3=(a+b)(a^2-ab+b^2)\]

OpenStudy (aravindg):

wat abt a^3-b^3

OpenStudy (aravindg):

and a^3+b^3+c^3-3abc

OpenStudy (anonymous):

\[a^3-b^3=(a-b)(a^2−ab+b^2)\]

OpenStudy (anonymous):

\[(a + b + c)(a^2 + b^2 + c62 − ab − bc − ca).\]

OpenStudy (aravindg):

wat??